OpenStudy (anonymous):

Evaluate e^(tA) where A =[2,3,-3; 4,2,-4; 4,3,-5] I have no idea how to start. If someone could explain the process that would be great.

OpenStudy (goformit100):

4,2,-4;

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

step1 diagonalize A by finding out the eigenvalues eigenvectors. Using characteristic equation. I got eigenvalues are: 2, -2, and -1. Calculate the eigenvectors correspond to those eigenvalue. You will have P , find P^-1 . use formula \[\huge\P^-AP = D\] to get D . That D is your matrix using to calculate e^At . the number in main diagonal line of D is the number of A in e^At for example. If (I say If because I don't calculate ) your D is \[\left[\begin{matrix}2 & 0&0\\ 0 & -2&0\\0&0&-1\end{matrix}\right]\]. when plug it into the formula \[\huge\e^At = P A P^-\] it will have the form something like \[e^At= P (\left[\begin{matrix}e^2t & 0&0 \\ 0 & e^-2t&0\\0&0&e^t\end{matrix}\right]P^-\] you still have 1 more step is calculate that e^At hope this helps. It's not easy or a piece of cake. It take long time to find out the answer. Fortunately, we have a path to do. You can do it, right? Good luck

OpenStudy (anonymous):

all e^(At), just typo. don't make mistake there. I mean don't think (e^A)* t